OpenStudy (anonymous):

12. The temperature was 56 degrees this morning but dropped one degree per hour through the rest of the day. Use the graph you made for the equation for this problem, to approximate the answer to the following question: What temperature was it 15 hours later, according to your graph? 13. The temperature was 56 degrees this morning, but dropped one degree per hour through the rest of the day. Read over the problem and look at the graph you made to decide how to answer this question. Would it make sense to predict the temperature in 45 hours? Why or why not?

OpenStudy (anonymous):

y=56-x so if x=15 y=41

OpenStudy (anonymous):

56 is your b in mx +b because it is constant, does not change in this instance, drops by 'one' thus variable x and no "m" as no coefficient

OpenStudy (anonymous):

56-x=45= 56-45=11 Would make sense to predict this in certain climates i.e. Mountains above 10,000 feet...but otherwise temperature change typically not that drastic but on a 24 hour cycle

OpenStudy (anonymous):

these are the questions & answers to the previous questions. 10. The temperature was 56 degrees this morning but dropped one degree per hour through the rest of the day. Write the linear equation for this problem in y=mx+b form. Note: this equation has a negative slope (m). Answer is: y= -x+56 11. The temperature was 56 degrees this morning but dropped one degree per hour through the rest of the day. Make a T-table for the equation for this problem, using the x-values 0, 5, 10. Type in the ordered pairs you get in the T-table as your answer. answer is: (0,56) (5,51) (10,46)